# DarkAI: Reconstructing the density, velocity and tidal field of dark matter from DESI-like bright galaxy sample

### v1

@@ 5 Summary and Conclusion
This work is supported by the National SKA Program of China (2022SKA0110200 and 2022SKA0110202), the National Key R&D Program of China (2023YFA1607800, 2023YFA1607801, 2023YFA1607804), the National Natural Science Foundation of China (Nos.12103037, 12261141691, 12373005), 111 project No. B20019, Shanghai Natural Science Foundation (Nos. 19ZR1466800, 24DX1400100, 23JC1410200) and Zhangjiang special support (No. ZJ2023-ZD-003). We acknowledge the science research grants from the China Manned Space Project (Grant No. CMS-CSST-2021-A02, CMS-CSST-2021-A03, CMS-CSST-2021-B01) and the Fundamental Research Funds for the Central Universities (Grant No. XJS221312). H.W. acknowledgements supports from the National Natural Science Foundation of China (NSFC, Nos. 12192224) and CAS Project for Young Scientists in Basic Research, Grant No. YSBR-062. The authors acknowledge the Beijing Super Cloud Center (BSCC) for providing HPC resources that have contributed to the research results reported in this paper. This work is also supported by the High-Performance Computing Platform of Xidian University.

### v2

@@ 8 Acknowledgments
This work is supported by the National SKA Program of China (2022SKA0110200 and 2022SKA0110202), the National Key R&D Program of China (2023YFA1607800, 2023YFA1607801, 2023YFA1607804), the National Natural Science Foundation of China (Nos.12103037, 12261141691, 12373005), 111 project No. B20019, Shanghai Natural Science Foundation (Nos. 19ZR1466800, 24DX1400100, 23JC1410200) and Zhangjiang special support (No. ZJ2023-ZD-003). We acknowledge the science research grants from the China Manned Space Project (Grant No. CMS-CSST-2021-A02, CMS-CSST-2021-A03, CMS-CSST-2021-B01) and the Fundamental Research Funds for the Central Universities (Grant No. ZYTS25248, XJS221312). This work is supported by the State Key Laboratory of Dark Matter Physics and the Young Data Scientist Program of the China National Astronomical Data Center (No.NADC2025YDS-01). H.W. acknowledges support from the National Natural Science Foundation of China (NSFC, Nos. 12192224) and CAS Project for Young Scientists in Basic Research, Grant No. YSBR-062. This work is supported by Natural Science Basic Research Program of Shaanxi (Program No. 2025JC-YBMS-016). The authors acknowledge the Beijing Super Cloud Center (BSCC) for providing HPC resources that have contributed to the research results reported in this paper. This work is also supported by the High-Performance Computing Platform of Xidian University.
Appendix A Testing on FA effect
To assess the potential impact of FA on the reconstruction, we simulate the FA effect using a four-pass observation strategy, consistent with the DESI survey design (Hahn et al., 2023; Bianchi et al., 2025). Each pass corresponds to a separate tiling of the survey footprint, targeting available galaxies while accounting for fiber collisions and survey geometry. After applying this simulation, the average completeness across the survey area is approximately 80%, reflecting the partial sampling introduced by the FA process. Figure 18 shows the resulting completeness map for the DESI footprint, highlighting variations in coverage due to the tiling strategy and the FA algorithm. This map illustrates the nonuniform spatial sampling that can introduce systematic biases in clustering measurements and reconstruction performance.
We then retrain the model using the sample with FA included. Figure 19 shows the ratios of the auto- and cross-correlation power spectra. The blue line represents results without the FA effect, while the red lines show results with the FA effect included. The green lines illustrate a simple correction using the nearest-neighbor galaxies. At large scales ($k\lesssim 0.1\,h\,\mathrm{Mpc}^{-1}$), all curves remain close to unity, indicating accurate reconstruction. However, at smaller scales, the FA effect introduces a noticeable suppression in the autopower spectrum (red lines), and the cross-power spectrum ratio also declines. The nearest-neighbor correction partially mitigates this suppression, as seen in the green lines. These results demonstrate that while the FA effect does not significantly impact LSSs, it biases small-scale clustering and highlights the need for more sophisticated corrections in future work. Nevertheless, we find that at small scales, the observed underestimation is dominated more by the limitations of the UNet prediction than by the FA effect itself.
Figure 18: Completeness map of the DESI survey footprint simulated using the four-pass observation strategy.
Figure 19: Ratios of the auto- (left panel) and cross-power (right panel) spectra between the reconstructed and true density fields, illustrating the impact of the FA effect. The blue lines show results without the FA effect and the red lines include the FA effect. The green lines represent a simple correction using nearest-neighbor galaxies.
Appendix B Comparison between Separately Trained and Unified Models
To assess the potential benefits of unified training, we conduct an additional experiment in which a single network is designed to reconstruct both the density and velocity fields simultaneously. This configuration contrasts with our original approach, where separate models were trained for each field. Due to hardware memory limitations, we performed this comparison using data cubes of size $256^{3}$ rather than the original $512^{3}$ cubes employed in our main analysis.
In Figure 20, we present a comparative analysis of the reconstruction performance between the separately trained models (blue lines) and the unified model (orange lines). Our results indicate that the overall reconstruction performance is largely consistent between the separate and unified training strategies, although noticeable discrepancies appear in certain metrics, such as the divergence and vorticity power spectra. In particular, the separately trained models demonstrate improved accuracy in reconstructing the divergence field, exhibiting lower scatter and reduced power suppression. Notably, these models still recover coherent structures across both the density and velocity fields, suggesting that feature consistency is well maintained even without joint training. This implies that despite the absence of shared parameters, separate training does not compromise the alignment between the reconstructed fields under the current experimental setup.
While it is generally expected that a unified model would produce more consistent features between the reconstructed density and velocity fields, the results indicate that under current memory constraints, training separate models remains the preferred strategy for maximizing accuracy across a wide range of scales. Future implementations on GPUs with larger memory capacity may revisit the unified approach as a viable and potentially more efficient alternative.
Figure 20: Comparison of the reconstruction performance between the separately trained models (blue lines) and the unified model (orange lines). As indicated in the panel titles, each subplot corresponds to one of the following diagnostic metrics: (1) the autopower spectrum of the dark matter field; (2) the cross-power spectrum between the reconstructed and true dark matter fields; (3) the momentum field power spectrum; (4) the velocity magnitude power spectrum, (5) the velocity divergence power spectrum; and (6) the vorticity power spectrum.